OpenStudy (anonymous):

A 1‐kg hockey puck at rest with Velcro wrapped around its side lays on top of an ice pond (consider this surface to be frictionless). A skater hits the puck and it moves with a constant speed of 8 m/s. The puck is headed towards another identical puck found at rest on the same pond. They collide, get stuck together, and move together as a single object from that point on. 1– When combined together, what is the weight (NOT MASS) of the moving pucks? 2 – What is the net force acting on the moving puck prior to the collision of pucks?

OpenStudy (anonymous):

mass and weight are equal on earth. so 2kg 8*1=8N

OpenStudy (anonymous):

There is more to it: 3 – What is the speed of the pucks when they move together after colliding? 4 – What is the momentum of the pucks when moving together? 5 – What is the kinetic energy of the combined pucks after they collide?

OpenStudy (anonymous):

I don't want the answer only. I want to know how to do these first

OpenStudy (anonymous):

4m/s and i cant help you with the rest unless you give me the equations. been over a year since i took physics

OpenStudy (anonymous):

that is all i was given

OpenStudy (anonymous):

umm.. 1 pucks are 1 kg, 2 pucks, 2*1=2 2. f=mass*velocity 8m/s * 1kg = 8N 3. 8N/2kg=4m/s

OpenStudy (anonymous):

yes. you ae supposed to know the formulas. they should be in your textbook. i need the mometum and kinetic energy formulas to solce the other problems

OpenStudy (anonymous):

oh, wait, i'm thinking of something else

OpenStudy (anonymous):

1. W = m*g 2. Force= mass x acceleration 3. momentum = mass*velocity 4. m1*u1 + m2*u2 = (m1+m2)*v 5. (1/2)*m*v^2

OpenStudy (anonymous):

found them. momentum = mass*velocity ke=mass*velocity/2

OpenStudy (anonymous):

momentum = 8 (4m/s *2kg) ke=2kg*4^2 /2 2*16/2 16

OpenStudy (anonymous):

i see, i never got the 4 in there. i didn't know you was to divide since i didn't know the velocity. i am horrible at this subject

OpenStudy (anonymous):

any good reference sites you know of that of help me?

OpenStudy (anonymous):

thank you for all the help by the way

OpenStudy (anonymous):

youre welcome. just google "physics help" or the type of problem to find a site

OpenStudy (anonymous):

i'll try that, plus, i see a book that i can read on amazon. thanks again.